OpenStudy (anonymous):

Given the quadratic function f(x) = 9x2 – 36x + 36, find a value of x such that f(x) = 9. (Points : 20)

OpenStudy (anonymous):

f(x) = 9x2 – 36x + 36 f(x) = 9 so 9x2 – 36x + 36 = 9 divide eqn with 9 we get, x^2 - 4x + 4 = 1 or x^2 - 4x + 3 = 0 x^2-3x-x+3=0 x(x-3)-1(x-3)=0 (x-1)(x-3)=0 so x-1=0 or x-3=0 so x=1 or x=3 so for x=1 and x=3, f(x)=9